OpenStudy (ladiesman217):

medal awarded Choose one of the factors of 27x3 + 512y3 a) 3 b)3x – 8y c)9x2 – 24xy + 64y2 d)9x2 + 24xy + 64y2

OpenStudy (ladiesman217):

i cant figure out how to do this

OpenStudy (ladiesman217):

can you help me hexagon001

OpenStudy (raden):

use the formula : a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (hexagon001):

im scratching my head too

OpenStudy (raden):

27x^3 + 512y^3 = (3x)^3 + (8y)^3

OpenStudy (raden):

now, use the formula above to get of factors a^3 + b^3 = (a+b)(a^2 - ab + b^2) we have (3x)^3 + (8y)^3 it means a = 3x and b = 8y therefore, (3x)^3 + (8y)^3 = (3x+8y)((3x)^2 - 3x(8y) + (8y)^2) (3x)^3 + (8y)^3 = (3x+8y)(9x^2 - 24xy + 64y^2)

OpenStudy (raden):

so, it is C @ladiesman217